OpenStudy (anonymous):

find the volume of the solid generated by rotating the region bounded by y= cos x, y=0, x=0, and x= pi/2 about y-axis

OpenStudy (anonymous):



OpenStudy (anonymous):

are you told to use a specific method?

OpenStudy (anonymous):

slice,disk, washer........

OpenStudy (anonymous):

I don't think you can get an exact answer.

OpenStudy (anonymous):

I think you can only set up an integral and approximate it by simpsons or trapeziodal rule.

OpenStudy (anonymous):

V= pi integral x^2 dy

OpenStudy (anonymous):

please help i really dont know this

OpenStudy (anonymous):

\[ V= \pi \int\limits_{0}^{1} (\cos^{-1} (y) )^2 dy\]

OpenStudy (anonymous):

You can't evaluate it exactly. All you can do is approximate it. Probably use simpsons rule with 5 function values, should give a pretty good answer.

OpenStudy (anonymous):

that;s if it was about the x axis.

OpenStudy (anonymous):

it says about y axis.

OpenStudy (jamesj):

sorry ... need my morning coffee

OpenStudy (anonymous):

maybe the asker made a typo.

OpenStudy (jamesj):

But even so, this integral is doable. Use integration by parts twice. First time with u' = 1, v = arccos^2(x). Remember that (d/dx)(arccos x) = -1/sqrt(1-x^2)

OpenStudy (anonymous):

how do you rotate the y=cos x about y-axis?

OpenStudy (anonymous):

i think its washer if you rotate it about y-axis

OpenStudy (whynot):

It is disk method if you rotate it about x or y axis